OpenStudy (vortish):

a pottery studio gives 3 and 5 hr classes for 25 dollars a hour and has collect 2200 for 24 classes how many of the 3 hr and how many 5 hours were paid for

OpenStudy (vortish):

any help would be appreciated nin I know you have helped me bf so thanks in advance

OpenStudy (vortish):

i dot know that why im asking for help if some one steers me in the right direction i can get it from there i just dont understand how to start it

OpenStudy (vortish):

some we have to get one variable on the left side of the equation so if we subtract x from both sides we get y=24-x correct?

OpenStudy (vortish):

yes it is a system of equation correct

OpenStudy (vortish):

mabey i kind of get frustrated with these kind of equations and i know the step just finding out which ones go were is were i screw it all up

OpenStudy (vortish):

i undestand i know to find x or y you have to get the variable on one side of the equation to solve for x or y

OpenStudy (vortish):

x =695 i think if i did it right 3x +5(24-x)=2200 3x +115 = 2200 divide both sides by three to get x = 695

OpenStudy (vortish):

if x = 695 and y = 23 im still lost

OpenStudy (aripotta):

you can't answer this...? unless they can sell half a class. i got 6.5 for y.

OpenStudy (aripotta):

am i doing something wrong here?

OpenStudy (aripotta):

did you write the question correctly?

OpenStudy (vortish):

i did 3(696) +5y=2200 the you get 2085 + 5y = 2200 subtract 2085 from both sides to get 5y = 115 then y = 23

OpenStudy (aripotta):

that can't be possible though, since they only had 24 classes total

OpenStudy (vortish):

yes i is correct ari

OpenStudy (aripotta):

unless they can sell half classes, this question cannot be possible :/ or maybe someone donated them money

OpenStudy (aripotta):

but still, it doesn't work

OpenStudy (aripotta):

i got y = 6.5?

OpenStudy (aripotta):

wat

OpenStudy (aripotta):

show me :o

OpenStudy (vortish):

ok that is were i goofed the pooch and other than that my math was right

OpenStudy (aripotta):

i don't get it :/

OpenStudy (aripotta):

i still don't understand :<

OpenStudy (aripotta):

but why 3 and 5?

OpenStudy (aripotta):

oh wait

OpenStudy (aripotta):

ok.

OpenStudy (aripotta):

oh.

OpenStudy (aripotta):

wouldn't i have been able to get the same answers without dividing by 25 though?

OpenStudy (aripotta):

maybe i just suck lol

OpenStudy (aripotta):

*back into corner crying*

OpenStudy (aripotta):

backs*

OpenStudy (aripotta):

no :/ i woke up at 1:30 today

OpenStudy (aripotta):

i'll try it again

OpenStudy (aripotta):

oooh i see what i did wrong

OpenStudy (aripotta):

i didn't combine -75y and 125y

OpenStudy (aripotta):

wait. yes i did :/ i don't know anything anymore

OpenStudy (aripotta):

i'll leave now :l

OpenStudy (aripotta):

i got it now. i multiplied 25 and 75, not 24 and 75